# SWE-bench Goes Live!

### v1

@@ Abstract
The issue-resolving task, where a model generates patches to fix real-world bugs, has emerged as a critical benchmark for evaluating the capabilities of large language models (LLMs). While SWE-bench and its variants have become standard in this domain, they suffer from key limitations: they have not been updated since their initial releases, cover a narrow set of repositories, and depend heavily on manual effort for instance construction and environment setup. These factors hinder scalability and introduce risks of overfitting and data contamination. In this work, we present SWE-bench-Live [^1] Homepage: https://swe-bench-live.github.io/, Code: https://github.com/SWE-bench-Live, Dataset: https://huggingface.co/SWE-bench-Live, a live-updatable benchmark designed to overcome these challenges. Our initial release consists of 1,319 tasks derived from real GitHub issues created since 2024, spanning 93 repositories. Each task is accompanied by a dedicated Docker image to ensure reproducible execution. Central to our benchmark is RepoLaunch, an automated curation pipeline that streamlines the entire process from instance creation to environment setup, removing manual bottlenecks and enabling scalability and continuous updates. We evaluate a range of state-of-the-art agent frameworks and LLMs on SWE-bench-Live, revealing a substantial performance gap compared to static benchmarks like SWE-bench, even under controlled evaluation conditions. To better understand this discrepancy, we perform detailed analyses across repository origin, issue recency, and task difficulty. By providing a fresh, diverse, and executable benchmark grounded in live repository activity, SWE-bench-Live facilitates rigorous, contamination-resistant evaluation of LLMs and agents in dynamic, real-world software development settings.
@@ 1 Introduction
    Staleness. SWE-bench and its derivatives have not been updated since their initial releases (Oct, 2023), making them static benchmarks. Because LLMs are trained on massive inscrutable corpora, these static datasets are at risk of data leakage, as they could have be been unpurposely included in model training data. This raises concerns about whether newer models are making truly generalizable progress or merely memorizing benchmark content, reducing the benchmarks’ effectiveness in distinguishing model capabilities.
@@
    Limited repository coverage. These benchmarks draw from a small set of 12 repositories, limiting diversity in codebases, domains, and programming practices (see Table 1 for details). This narrow scope weakens the generalizability and robustness of evaluations.
@@
To address these challenges, we introduce SWE-bench-Live, a live, automated, and scalable benchmark designed to evaluate LLMs on real-world issue resolution tasks. Unlike recent efforts such as LiveCodeBench jain2024livecodebench , which focus on algorithmic programming problems, SWE-bench-Live is the first continuously updating benchmark tailored for complex, repository-level tasks that require multi-file reasoning, environment setup, and reproducible execution.
Figure 1 presents the construction pipeline of SWE-bench-Live. At the heart of this framework lies RepoLaunch, a fully automated system that eliminates manual bottlenecks by streamlining the entire process—from issue mining to environment packaging. Specifically, RepoLaunch employs an agentic, end-to-end workflow to set up Docker-based environments by identifying relevant instruction files, selecting appropriate base images, installing dependencies, building the project, and validating its test suite. This automation enables continuous updates, broad repository coverage, and scalable dataset expansion. The current release of SWE-bench-Live includes 1,319 issue-resolution tasks sourced from real-world GitHub issues created since 2024, spanning 93 repositories. Compared to existing benchmarks, this marks a significant advancement in freshness, diversity, and scale (see Table 1).
We evaluate three leading agent frameworks (i.e., OpenHands openhands , SWE-Agent sweagent , and Agentless xia2024agentless ) in combination with four state-of-the-art LLMs (namely, GPT-4.1, GPT-4o, Claude 3.7 Sonnet, and DeepSeek V3). Our results reveal a substantial performance gap between SWE-bench-Live and prior static benchmarks. For instance, the best-performing agent–model combination, OpenHands with Claude 3.7 Sonnet, achieves a resolved rate of only 19.25% on SWE-bench-Live. Even under controlled conditions with identical evaluation protocols, the same setup yields a resolved rate of 43.20% on SWE-bench Verified, more than double its performance on SWE-bench-Live. Further breakdowns by repository origin and instance difficulty suggest that this discrepancy stems not only from benchmark familiarity but also from the greater diversity of SWE-bench-Live. These findings highlight the limitations of static, manually curated benchmarks and underscore the importance of dynamic, automatically updated testbeds for advancing robust and generalizable code agent systems.
@@
Together, these contributions establish a new and robust standard for evaluating the true capabilities of code LLMs and agent-based systems.
@@ 3 SWE-bench-Live
Targeting the issue resolution task on real-world GitHub repositories, SWE-bench serves as a practical proxy for evaluating the coding capabilities of LLM-based systems. The issue resolving task is defined as follows: given a code repository and an associated issue, an approach (e.g., LLM agent) is required to generate a patch that resolves the issue and passes the test cases.
@@
Thanks to its fully automated pipeline, SWE-bench-Live can be maintained with minimal, ideally zero manual effort. We plan to update SWE-bench-Live on a monthly basis, continually providing the community with an up-to-date evaluation dataset. This enables contamination-free, rigorous assessment of AI systems’ issue-resolving capabilities in a constantly evolving real-world setting.
3.1 Task Formulation
Issue resolving is the task that introduced by SWE-bench swebench for benchmarking AI coding capabilities. In simple terms, it simulates the process of a developer submitting a pull request to address an issue. The formulation of the issue-resolving task is illustrated in Figure 2.
Generating Patch.
Evaluating Patch.
@@ 3.2 Raw Issue–PR Crawling
The first phase of the SWE-bench-Live pipeline involves collecting real-world issue–PR pairs from popular open-source GitHub repositories.
@@
We focus on Python repositories for the initial release of SWE-bench-Live, aligning with SWE-bench and other prior benchmarks, as Python is one of the most widely used languages in open-source development and well-supported by existing LLMs.
The selection process consists of three filtering stages, designed to ensure both the quality and relevance of benchmark data:
  • We first queried the GitHub API for repositories with over 1,000 stars and Python set as the primary language. The star threshold helps identify mature and actively maintained projects with a large user base, while language filtering ensures consistency in model evaluation. This initial query yielded 8,577 repositories as of April 2025.
  • To further ensure repository richness and task availability, we refined this set by requiring each repository to have more than 200 combined issues and pull requests, over 200 forks (as a proxy for community engagement and adoption), and at least 60% of its codebase written in Python (to avoid polyglot or non-representative projects). This reduced the pool to 3,316 repositories.
  • Finally, to comply with legal and ethical usage guidelines, we retained only repositories containing a valid open-source license, resulting in a final selection of 2,609 repositories suitable for benchmark construction.
@@ 3.3 RepoLaunch: Automated Execution Environment Setup
A critical requirement for test-based evaluation in issue-resolution benchmarks is the availability of a faithful execution environment—one that reproduces the exact conditions under which a bug was introduced. While issue–PR pairs provide textual and code-level artifacts, without an execution environment, these cannot be tested or validated in a reproducible manner.
However, constructing such environments remains one of the most challenging and labor-intensive bottlenecks in dataset creation. Prior work such as SWE-bench swebench and SWE-Gym swegym rely entirely on manual setup, with SWE-Gym reporting over 200 hours of effort for a relatively small dataset. This challenge is exacerbated by three key factors: (i) lack of standard build instructions across repositories, (ii) version drift and dependency incompatibilities due to time decay, and (iii) the need for fine-grained environments at the snapshot level, since different commits may require different setups even within the same repository.
To address this scalability bottleneck, we introduce RepoLaunch, an LLM-powered agentic framework that automates the end-to-end construction of Dockerized execution environments for arbitrary repository snapshots. Unlike prior approaches that rely on scripted heuristics or manual curation, RepoLaunch simulates the trial-and-error reasoning process of human developers using an interactive, ReAct-style agent yao2023react ; zhang2024ufo .
@@
Each benchmark instance is tied to a repository snapshot defined by the base commit of an issue. An execution environment is considered valid if (i) the project builds from source and (ii) its test suite passes with zero or acceptable failures. This environment serves as the ground truth for verifying the correctness of model-generated patches.
Automated Agentic Workflow.
At the core of RepoLaunch lies an agentic workflow that emulates how human developers approach unfamiliar codebases: reading documentation, iteratively issuing commands, debugging errors, and refining their strategy based on feedback. Unlike static scripts or prompt-only approaches, our design instantiates a multi-role, interactive agent framework grounded in the ReAct paradigm yao2023react ; zhang2024ufo . This workflow is composed of two decoupled but coordinated agents: a setup agent that incrementally builds the environment, and a verification agent that checks for correctness. This separation of concerns provides modularity, simplifies control logic, and enables scalable error recovery.As shown in Figure 1, RepoLaunch proceeds through five key stages:
  • Relevant Files Identification. The first step is to identify relevant files in the repository–such as CI/CD pipelines and README files that are likely to contain useful information for setting up the environment (a detailed list is provided in the Appendix E).
@@
  • Verification Agent. Once the setup agent signals readiness or hits a step limit, a separate verification agent takes over. Its role is to determine the appropriate test command (e.g., pytest, tox, or a custom script), execute it, and evaluate results. If test failures are observed, the feedback is passed back to the setup agent, which re-enters the ReAct loop to repair the environment. This design enables retryable, stateful, and verifiable setup.
  • Finalization. Once all tests pass, the resulting container is committed as a snapshot-specific Docker image. This ensures reproducibility and enables downstream usage for model evaluation.
The agentic workflow in RepoLaunch is a principled and modular design that goes beyond simple command execution. It separates setup and verification logic for maintainability, supports iterative self-correction to handle noisy or incomplete metadata, and mirrors real developer behavior for greater robustness. Combined with time-aware dependency control and snapshot-level granularity, this design enables the first fully automated, scalable benchmark pipeline for real-world issue resolution.
Time Machine for Version Incompatibility Resolution.
A significant challenge in setting up environments for historical repository snapshots is dependency version drift, where unpinned dependencies are resolved to the latest package versions, leading to backward incompatibilities that frequently break the build. This is particularly problematic for older codebases that were never tested against modern package ecosystems. To address this, we introduce a time-machine mechanism that constrains the dependency resolver to only consider package versions released prior to the snapshot’s base commit timestamp. Concretely, we modify the default pip index server to route through a custom proxy that filters packages by release date. This simple yet effective solution mitigates the risk of “future” incompatibilities and substantially improves environment setup success rates across out-of-date repositories.
We open-source RepoLaunch  [^3] https://github.com/microsoft/SWE-bench-Live to benefit the broader community. While originally developed to automate the most labor-intensive stage of benchmark construction, namely environment setup, RepoLaunch significantly reduces the need for manual intervention, enabling scalable and reproducible benchmarking at scale. Beyond this primary use case, RepoLaunch also serves as a practical tool for developers, offering a fast and reliable way to set up execution environments for unfamiliar or legacy codebases. Its ability to reconstruct historical setups and resolve dependencies automatically makes it broadly useful in both research and real-world development workflows.
@@ 3.4 Validating Task Instances
To ensure the quality of the benchmark, each task instance is further validated to confirm that the associated PR effectively resolves the issue it is intended to fix. The validation is based on analyzing changes in the test suite results before and after applying the PR’s patch. Specifically, we focuses on identifying two key behaviors in the test outcomes:
@@ 3.5 SWE-bench-Live Statistics
To support lightweight experimentation, we construct a lite subset of SWE-bench-Live, referred to as SWE-bench-Live-Lite, by sampling 50 instances per month from issues created between October 2024 and March 2025. This results in a compact set of 300 instances that balances recency, diversity, and evaluation efficiency.
@@
Figure 4: Repository distribution.
Table 2: Statistics of SWE-bench-Live. *: Python code only. : Statistics of gold patches.
@@ 4.1 Setups
Agents and Models.
To evaluate the effectiveness of our proposed SWE-bench-Live, we conduct experiments using three representative agent frameworks. These include the general-purpose coding agent OpenHands openhands (paired with the CodeAct agent wang2024executable ), as well as two agents specifically designed for issue-resolving tasks: SWE-Agent sweagent and Agentless xia2024agentless . For OpenHands, we set a maximum of 60 iterations per instance. For SWE-Agent, we limit the number of LLM calls to 100 per instance to maintain computational efficiency. For Agentless, we largely follow the original pipeline, which consists of two main stages: issue localization and patch generation. However, we omit the reranking stage based on regression testing, as supporting this step on SWE-bench-Live would require substantial infrastructure adaptation and is beyond the scope of this study. Consequently, both the localization and repair stages in our Agentless evaluation produce a single sample without reranking. We test these agents using four recent state-of-the-art LLMs, covering both proprietary and open-source models: GPT-4o openai_gpt4o (gpt-4o-2024-11-20), GPT-4.1 openai_gpt41 (gpt-4.1-2025-04-14), Claude 3.7 Sonnet claude_37 (claude-3-7-sonnet-20250219), and DeepSeek V3 deepseek_v3 (DeepSeek-V3-0324).
Table 3: Performance on SWE-bench-Live-Lite.

### v2

@@ Abstract
The issue-resolving task, where a model generates patches to fix real-world bugs, has emerged as a key benchmark for evaluating the capabilities of large language models (LLMs). While SWE-bench has become the dominant benchmark in this domain, it suffers from several limitations: it has not been updated since its release, is restricted to only 12 repositories, and relies heavily on manual effort for constructing test instances and setting up executable environments, significantly limiting its scalability. We present SWE-bench-Live [^1] Homepage of SWE-bench-Live: https://swe-bench-live.github.io/, Code at https://github.com/SWE-bench-Live, and Dataset at https://huggingface.co/SWE-bench-Live, a live-updatable benchmark designed to address these limitations. Our initial release includes 1,319 tasks derived from real GitHub issues created since 2024, spanning 93 repositories. Each task is accompanied by a dedicated Docker image to ensure reproducible execution. Additionally, we introduce an automated curation pipeline that streamlines the entire process from instance creation to environment setup, removing manual bottlenecks and enabling scalability and continuous updates. We evaluate a range of state-of-the-art models and agent frameworks on SWE-bench-Live, offering detailed empirical insights into their real-world bug-fixing capabilities. By providing a fresh, diverse, and executable benchmark grounded in live repository activity, SWE-bench-Live supports reliable, large-scale assessment of code LLMs and code agents in realistic development settings.
@@ 1 Introduction
    Staleness. SWE-bench and its derivatives have not been updated since their initial releases, making them static benchmarks. Because LLMs are trained on massive inscrutable corpora, these static datasets are at risk of data contamination, as they could have be been unpurposely included in model training data. This raises concerns about whether newer models are making truly generalizable progress or merely memorizing benchmark content, reducing the benchmarks’ effectiveness in distinguishing model capabilities.
@@
    Limited repository coverage. These benchmarks draw from a small set of repositories, limiting diversity in codebases, domains, and programming practices (see Table 1 for details). This narrow scope weakens the generalizability and robustness of evaluations.
@@
To address these challenges, we introduce SWE-bench-Live, a live and scalable benchmark built for evaluating LLMs on real-world issue resolution tasks. In contrast to recent efforts such as LiveCodeBench jain2024livecodebench , which target algorithmic programming problems, SWE-bench-Live is the first live-updating benchmark designed for complex, repository-level tasks that demand multi-file reasoning, environment setup, and reproducible execution. Figure 1 illustrates the construction pipeline of SWE-bench-Live. At the core of our framework is RepoLaunch, a fully automated pipeline that eliminates manual bottlenecks by streamlining the entire process—from issue mining to environment packaging. More specifically, RepoLaunch leverages an agentic and end-to-end workflow to setup the Docker environment by identifying relevant instruction files, selecting base images, installing necessary dependencies, building the project, and validating its test suite. This automation enables continuous updates, broad repository coverage, and large-scale dataset expansion. Our current release of SWE-bench-Live contains 1,319 issue-resolution tasks sourced from real-world GitHub issues created since 2024, spanning 93 repositories. Compared to existing benchmarks, this represents a significant leap in freshness, diversity, and scale (see Table 1).
We evaluate three leading agent frameworks (i.e., OpenHands openhands , SWE-Agent sweagent , and Agentless xia2024agentless ) in combination with four state-of-the-art LLMs (namely, GPT-4.1, GPT-4o, Claude 3.7 Sonnet, and DeepSeek V3). Consistent with performance rankings reported on SWE-bench Verified, [^3] https://www.swebench.com/ we observe that OpenHands, when paired with Claude 3.7 Sonnet, achieves the highest performance on SWE-bench-Live. However, its overall results are significantly lower compared to those achieved on SWE-bench Verified. To explore this discrepancy further, we conduct a controlled comparison and find that the same agent-LLM pair consistently performs worse on SWE-bench-Live than on SWE-bench. This finding suggests that existing models may be overfitting to static benchmarks like SWE-bench, underscoring the importance of developing more dynamic and diverse evaluation settings, such as those provided by SWE-bench-Live.
@@ 3 SWE-bench-Live
Targeting the issue resolution task on real-world GitHub repositories, SWE-bench serves as a practical proxy for evaluating the coding capabilities of LLM-based systems. The issue resolving task is defined as follows: given a code repository and an associated issue, an approach (e.g., LLM agent) is required to generate a patch that resolves the issue and passes the test cases (see Appendix B for details).
@@
Thanks to its fully automated pipeline, SWE-bench-Live can be maintained with minimal–ideally zero–manual effort. We plan to update SWE-bench-Live on a monthly basis, continually providing the community with an up-to-date evaluation dataset. This enables contamination-free, rigorous assessment of AI systems’ issue-resolving capabilities in a constantly evolving real-world setting.
@@ 3.1 Raw Issue–PR Crawling
The first phase of the SWE-bench-Live pipeline involves collecting real-world issue–pull request (PR) pairs from popular open-source GitHub repositories.
@@
We focus on Python repositories for the initial release of SWE-bench-Live, aligning with SWE-bench and other prior benchmarks due to its popularity. The selection process includes three filtering stages: (i) We first queried GitHub API for repositories with over 1,000 stars and Python set as the primary language. This initial query yielded 8,577 repositories as of April 2025. (ii) We then refined this set by requiring each repository to have more than 200 issues and pull requests, over 200 forks, and at least 60% of its codebase written in Python. This reduced the pool to 3,316 repositories. (iii) Finally, to comply with licensing requirements, we retained only repositories containing a valid open-source license, resulting in a final selection of 2,609 repositories.
@@ 3.2 RepoLaunch: Automated Execution Environment Setup
The “raw” issue–PR pairs remain at the textual and plain code level. To support subsequent test-based evaluation, it is required to provide an execution environment capable of running tests locally and producing execution feedback. In the context of issue-resolving benchmarks, the execution environment is critical for test-based evaluation.
However, preparing such execution environments is widely recognized as the most labour-intensive step in constructing issue-resolving datasets. In prior work, including SWE-bench swebench and SWE-Gym swegym , environment setup has been performed entirely by humans. For example, SWE-Gym reports that building execution environments required over 200 hours of manual effort, underscoring a significant scalability bottleneck. Notably, even repository-level environments are insufficient: different commits within the same repository may depend on different libraries or configurations, necessitating environment construction at the snapshot level. SWE-bench partially mitigates this by building environments per version tag, but the granularity remains coarse and relies on manual labor.
To address this bottleneck, we introduce an agent-based framework RepoLaunch, which automatically creates a fully functional execution environment for each issue instance. For any given repository snapshot, RepoLaunch produces a Docker container that installs all required dependencies, builds the project, and validates its test suite. This containerized instance serves as the foundation for running and evaluating model-generated patches.
@@
A repository snapshot corresponds to the codebase at the base commit associated with an issue. The goal is to recreate an environment faithful to that moment in time. We define a valid execution environment as a Docker container where (i) the codebase is correctly installed from source, and (ii) the repository’s test suite passes with zero or tolerable failures. This environment is essential for test-based evaluation, providing the ground truth mechanism to verify whether the issue has been resolved.
RepoLaunch follows an LLM-driven, agentic workflow zhang2024large ; zhang2025api inspired by how human developers set up unfamiliar projects, as shown in Figure 1. The process proceeds in five steps:
  • Relevant Files Identification. The first step is to identify relevant files in the repository–such as CI/CD pipelines and README files that are likely to contain useful information for setting up the environment (a detailed list is provided in the Appendix G).
@@
  • Verification. Once the setup agent determines that the environment has reached a satisfactory state or a step limit is reached, control is transferred to a verifying agent. The agent attempts to generate the appropriate test command and execute it. The execution results are evaluated with the agent to check if all test cases passed. If test failures occur, the results are fed back to the setup agent for further refinement. If all tests pass, the environment is considered valid.
  • Finalization. Upon successful validation, the container is committed as a Docker image, producing a instance-level execution environment for reuse.
Challenges of Version Incompatibility.
A major challenge when setting up out-of-date repositories is the “dependency version drift” issue. When dependencies are not pinned to specific versions, tools like pip by default will resolve to the latest package versions, which often introduce backward-incompatible issues and make the environment setup fail. To address this, we implement an time-machine mechanism by forcing the package installation tool to only look at valid versions released no later than the current base commit timestamp. Specifically, we modified the pip default index server to a proxy which fetches those valid package versions. This simple but effective strategy prevents the “future” version incompatibilities and significantly improves setup success rates.
We will open-source RepoLaunch to benefit the community. While designed for automated benchmark construction, RepoLaunch can also assist developers in quickly setting up environments for unfamiliar codebases. Its ability to replicate historical setups and automatically resolve environment dependencies positions it as a practical tool with broader applicability beyond benchmarking.
@@ 3.3 Validating Task Instances
To ensure the quality of the benchmark, each task instance is validated to confirm that the associated PR effectively resolves the issue it is intended to fix. The validation is based on analyzing changes in the test suite results before and after applying the PR’s patch. Specifically, we focuses on identifying two key behaviors in the test outcomes:
@@ 3.4 SWE-bench-Live Statistics
To support lightweight experimentation, we construct a lite subset of SWE-bench-Live by sampling 50 instances per month from issues created between October 2024 and March 2025. This results in a compact set of 300 instances that balances recency, diversity, and evaluation efficiency.
@@
Figure 3: Repository classifications.
Table 2: Statistics of SWE-bench-Live
@@
*Only count Python code. Stats of gold patch.
@@ 4.1 Setups
Agents and Model Selection.
To evaluate the effectiveness of our proposed SWE-bench-Live, we conduct experiments using three representative agent frameworks. These include the general-purpose coding agent OpenHands openhands (paired with CodeAct), as well as two agents specifically designed for issue-resolving tasks: SWE-Agent sweagent and Agentless xia2024agentless . For OpenHands, we set a maximum of 60 iterations per instance. For SWE-Agent, we limit the number of LLM calls to 100 per instance to maintain computational efficiency. For Agentless, we largely follow the original pipeline, which consists of two main stages: issue localization and patch generation. However, we omit the reranking stage based on regression testing, as supporting this step on SWE-bench-Live would require substantial infrastructure adaptation and is beyond the scope of this study. Consequently, both the localization and repair stages in our Agentless evaluation produce a single sample without reranking. We test these agents using four recent state-of-the-art LLMs, covering both proprietary and open-source models: GPT-4o openai_gpt4o (gpt-4o-2024-11-20), GPT-4.1 openai_gpt41 (gpt-4.1-2025-04-14), Claude 3.7 Sonnet claude_37 (claude-3-7-sonnet-20250219), and DeepSeek V3 deepseek_v3 (DeepSeek-V3-0324).
Evaluation Metrics.
Following the evaluation protocol of SWE-bench swebench , we adopt the Resolved Rate (%) as our primary metric. This measures the proportion of issues successfully resolved by the agent across all task instances. We also report the Patch Apply Rate (%), which indicates the percentage of generated patches that are syntactically correct and can be successfully applied to the codebase without errors. Additionally, we measure the Localization Success Rate (%) at the file level. This reflects whether the set of files modified by the generated patch matches the gold patch.
4.2 Performance on SWE-bench-Live
We report the performance of all agent–model combinations on the Lite subset of SWE-bench-Live in Table 3. Meanwhile, Table 4 presents the results of the top three combinations selected based on Lite performance, evaluated on the full version of SWE-bench-Live.
We observe that the same methods achieve substantially higher scores on SWE-bench compared to their performance on SWE-bench-Live, despite both benchmarks targeting the same issue-resolving task with identical settings. For example, recent state-of-the-art agents and models report a resolved rate exceeding 60% on the SWE-bench Verified subset [^4] https://www.swebench.com/. In contrast, the highest resolved rate on SWE-bench-Live is only 19.25%. Considering that the experimental setups on the SWE-bench leaderboard often involve dramatically high rollout numbers or iteration efforts, we specifically re-ran the best performing combination, OpenHands with Claude 3.7 Sonnet, on the SWE-bench verified subset using the exact same setups as in our experiments. The resulting resolved rate reached 43.20%, more than twice the score achieved on SWE-bench-Live. This is a particularly interesting phenomenon, as it highlights the challenges of constructing a benchmark that can objectively measure an AI system’s ability to resolve arbitrary and previously unseen issues. It also raises concerns about potential overfitting to SWE-bench. Similar phenomena are also observed in other existing issue-resolving datasets: the best-performing method in Multi-SWE-bench achieves a resolved rate of only 19.32%, while the highest score reported in OmniGIRL is as low as 8.6%.
Table 3: Performance on SWE-bench-Live (Lite subset).